OpenStudy (anonymous):

(k+x)(1+3x)^8 k is a constant,when coefficient of x=0 find the value of k

OpenStudy (anonymous):

Please reformulate your question. I don't get it.

OpenStudy (kc_kennylau):

When constant in (1+3x)^8 is obviously 1, and the coefficient of x in (1+3x)^8 is 24x (ask me if you don't know how I got this)

OpenStudy (kc_kennylau):

Therefore the coefficient of x in (k+x)(1+3x)^8 is... hmm... I think you know how to calculate it now don't you

OpenStudy (anonymous):

Just plug in x=0 in the equation.

OpenStudy (anonymous):

Can you do that?

OpenStudy (kc_kennylau):

And @JoelSjogren the question is comprehensible enough

OpenStudy (anonymous):

I disagree but I get it thanks to you.

OpenStudy (kc_kennylau):

Oh okay then

OpenStudy (anonymous):

Isn't x going to be 0 if its coefficient is 0x = 0.

OpenStudy (anonymous):

@kc_kennylau

ganeshie8 (ganeshie8):

x=0 wud make all other terms also 0, except the constant term.. so....

OpenStudy (anonymous):

the value of k is - 1/24 ?

OpenStudy (kc_kennylau):

exactly

OpenStudy (kc_kennylau):

I think so

ganeshie8 (ganeshie8):

(k+x)(3x+1)^8 (k+x)[(8 0)(3x)^8*1^0 + (8 1) (3x)^7 * 1^1 + .... (8 7)(3x)^1 1^7 + (8 8) (3x)^0 1^8 ]

OpenStudy (kc_kennylau):

@ganeshie8 please don't expand it fully...

ganeshie8 (ganeshie8):

you want to consider last two terms @AnImEfReaK those give u coefficients of x

ganeshie8 (ganeshie8):

??

OpenStudy (anonymous):

Yes i know now.

OpenStudy (anonymous):

the value of K? @ganeshie8

ganeshie8 (ganeshie8):

-1/24 is right

OpenStudy (anonymous):

thanks..

OpenStudy (anonymous):

As a side note, you can also do this using calculus. Let the expression be y and equate the constant term of y' with 0.

OpenStudy (kc_kennylau):

Wow I never thought of it

OpenStudy (anonymous):

@JoelSjogren I mean whats the use of taking the derivative?

OpenStudy (anonymous):

\[y = (k + x)(1 + 3x)^8\] By the product rule and the chain rule, \[y \prime = 1(1 + 3x)^8 + (k + 3x)8(1 + 3x)^7 3\] We want the constant term, so we look for the value when x = 0: \[y' = 1 + k*24\] This is 0. Solving for k gives k = -1/24

OpenStudy (anonymous):

Boblovesmath, it's beautiful. Other than that, I don't know the use of it.

OpenStudy (anonymous):

@JoelSjogren I know but I mean why did you take the slope of the tangent line at that point? Why did you do the derivative?

OpenStudy (anonymous):

The slope of the tangent line is the coefficient of x.

OpenStudy (anonymous):

@JoelSjogren Ohhhh!!!!! I got it now thanks!!! Thanks for putting it forward, I really appreciate it.

OpenStudy (anonymous):

No problem!

OpenStudy (loser66):

@JoelSjogren Beautiful job!!! never know this way. Thanks for posting.